W(t) is in Quadrant I with sin(t) = 2sqrt(2)/3. Find cos(t).
Answer by lwsshak3(11628)

W(t) is in Quadrant I with sin(t) = 2sqrt(2)/3. Find cos(t).
**
2 ways you can do it:
Using trig identity; sin^2+cos^=1
or the Pythagorean Theorem
I will do the 2nd method, leaving you to do the first method
..
working with a right triangle with the opposite side=2√2 and hypotenuse=3
adjacent side=√[3^2-(2√2)^2]=√(9-8)=√1=1
cos(t)=adjacent/hypotenuse=1/3